On aircraft carriers, catapults are used to accelerate jet aircraft to flight speeds in a short distance. One such catapult takes a 18,200-kg jet from 0 to 65 m/s in 2.9 s. (Assume the catapult acts in the positive horizontal direction. Indicate the direction with the sign of your answer where appropriate.)

(a) What is the acceleration of the jet in m/s2?

(b) What is the acceleration of the jet in g's?

(c) How far does the jet travel while it is accelerating?

(d) How large is the force that the catapult must exert on the jet?
Answer #1

a) 65/2.9 = 22.41 m/s^2

b) a (in g’s) = 22.41/9.8 = 2.29 g’s

b)d = 1/2 a t^2 = 0.5 x 22.41 x 2.9^2 = 94.25 m

c)F = ma = 18,200 x 22.41 = 407,862 N.